2 trains leave the station at the same time, one heading east and the other west. The eastbound train travels 10 mph faster. 3 hours later the trains are 510 miles apart. What it the rate of the eastbound train?

Answer by mananth(16949)

Distance = 510 miles
west bound train speed =x
east bound train = x+10
they are moving away.
so combined speed = x+x+10=2x+10
Time = 3hours
Distance = speed * time
510=(2x+10)*3
510/3 = 2x+10
170=2x+10
170-10=2x
160=2x
x= 80 mph
East bound train speed = 80+10 = 90 mph
m.ananth@hotmail.ca